# Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects

### v2

@@ 5.3 Instruction-tuning Effects
We report the percentage gains in rank-biserial correlation for the instruction-tuned variants over their base variants across the four datasets in Table 3. To avoid confusion, these results examine the effect of extracting surprisals from an instruction-tuned variant over a base variant with the same experimental setting. We do not add a prompt/instruction to the instruction-tuned variants inputs. The results show that instruction-tuning does not always improve the correlation between surprisal and novelty annotations. Only Llama instruction-tuned variants could improve the correlations on the VUA datasets; otherwise, instruction-tuning deteriorates the correlations.
@@ 5.4 Cloze-surprisal
In Table 4, we report the percentage gains in rank-biserial correlations for the cloze-surprisal values over the direct-surprisal values from the same model variants across the four datasets. We find cloze-surprisal to be boosting the correlations with extra points in many cases; however, it also deteriorates the correlations heavily in other cases. In Figure 3, we can observe the distribution of cloze-surprisal of the metaphoric words shifted away from zero and approaching a normal distribution. Also in Figure 4, we observe the same effect. However, against our intuition, the effect of cloze-surprisal is not more impactful in cases where the metaphor is located earlier in the sentence than later. The effect of cloze-surprisal, although positive, is not related to the position of the metaphoric word in the sentence.
For the GPT2 model family, cloze-surprisal consistently improves the corpus-based datasets’ correlations, achieving the strongest correlations in this study (out of the 32 recorded surprisals on each dataset) coming from GPT2-base on VUA-ratings with values of $r=.499$, $\rho=.499$, $r_{b}=.687$ and $auc=.843$. However, it deteriorates the correlations at large scales on the synthetic datasets.
Llama and Qwen model families also show positive gains in most of the cases. Most importantly, unlike the GPT2 model family, it boosts correlations on the synthetic datasets except for a few cases. Cloze-surprisal achieves the overall strongest correlation on the Lai2009 dataset using the Qwen-14B model, with $r_{b}=.539$ and $auc=.779$. Also, it achieves the overall strongest correlation on the GPT-4o-metaphors dataset using the Llama-8B model, with $r_{b}=.684$ and $auc=.842$.
@@
ppl. Nov. % $r_{b}$ Nov. % $r_{b}$
Fiction 108 2.97 .693 3.56 .478
News 89 2.80 .653 3.88 .413
Academic 73 1.85 .588 1.85 .413
Conversation 134 1.41 .482 0.62 .800
All 96 2.33 .638 2.70 .581
@@ 6 Discussion
Across multiple novelty annotation setups—continuous human ratings, dictionary-based binary labels, linguist-designed contrasts, and LLM-generated conventional vs. novel senses—we find consistently moderate associations between LM surprisal and metaphor novelty (best: $r=.49,\rho=.50,r_{b}=.69,\mathrm{AUC}=.84$). While these results are not directly comparable to prior work due to differences in datasets and task formulations, their magnitudes are broadly in line with reported surprisal–behavior associations in acceptability judgments Tjuatja et al. (2025) and reading times Oh and Schuler (2023a). The closest point of reference in the metaphor novelty literature is Parde and Nielsen (2018), who report $r=.44$ when predicting novelty scores from a wide range of linguistic features. They additionally report moderate inter-annotator agreement, underscoring that novelty itself is challenging to measure reliably by humans.
At the same time, surprisal has clear theoretical limits as a standalone predictor: interpreting a metaphor involves semantic integration and cross-domain mapping, not only predictability. We therefore expect surprisal to be most informative when combined with measures that more directly target novel interpretations and domain mappings, especially for highly creative (novel) metaphors.
@@
The negative effect of model size observed in corpus-based datasets mirrors the inverse scaling effect reported in reading time studies Oh and Schuler (2023b); Wilcox et al. (2025). Recently, Oh et al. (2024) argued that this effect is largely driven by word frequency, with larger models assigning increasingly non-human-like expectations to rare words. Since corpus-based novelty scores are often correlated with word frequency Do Dinh et al. (2018); Reimann and Scheffler (2024), frequency can be an underlying factor for this negative scaling effect.
By contrast, the synthetic datasets control for lexical identity and frequency: conventional and novel senses are elicited for the same set of words, and stimuli are constructed to cleanly separate the two senses. Under these controls, scaling improves alignment with novelty labels, agreeing with reports of positive scaling effects in other behavioural settings such as acceptability judgments and other reading time studies Tjuatja et al. (2025); Wilcox et al. (2023).
Overall, we suggest that these diverging scaling effects are due to the nature of the dataset types. While corpus-based datasets reflect metaphor novelty mainly through lexical properties such as word frequency, synthetic datasets more directly isolate the conventional–novel distinction by controlling for lexical properties confounds.
@@ 7 Conclusion
We acknowledge several limitations in this work. First, the scarcity of high-quality metaphor novelty annotations in existing literature constrains both coverage and generalizability. Second, we rely on pretrained language models whose training data and processes are not fully disclosed.
Finally, although we carefully analyse the effects of model architecture, size, and domain (genre), future work could adopt mixed-effects models to test the interaction of these variables.
@@ Acknowledgments
We acknowledge the reviewers for their valuable comments and feedback. And we also thank Vicky Tzuyin Lai and Tiago Pimentel for their responsiveness to our inquiries.
Appendix A Details of the datasets
@@ A.1 Statistics and Examples
5. One Mr Clarke can not duck away from if he wants to avoid a second Winter of Discontent .. -0.094 conventional 3.165 7.454
@@ A.2 GPT-4o-metaphors Construction
Appendix B Visualisations
B.1 Distributions
@@
Figure 3: Histograms of metaphor novelty, and direct surprisal (dark colour) of content metaphoric words (15,155) across 10 LMs. Additionally, cloze-surprisal of the same models is plotted in light colour, and surprisal from instruct-tuned variants (whenever applicable) is plotted in lined solid colour.
Figure 4: Scatter plots between metaphor novelty scores of VUA-ratings on the x-axis and surprisal (from GPT2-base) direct (blue) and cloze (red) on the y-axis at different relative positions of metaphoric words in sentences. 0-10% means metaphoric words located at the first 1/10Length1/10*Length of the sentence.

### v3

@@ 5.3 Instruction-tuning Effects
We report percentage gains in rank-biserial correlation for the instruction-tuned variants over their base variants across the four datasets in Table 3. This shows the effect of extracting surprisals from an instruction-tuned variant over a base variant with the same experimental setting. We do not add a prompt/instruction to the instruction-tuned models’ inputs. The results suggest that instruction-tuning does not always improve the correlation between surprisal and novelty annotations. Only Llama instruction-tuned variants could improve the correlations over their base variants on the VUA datasets; otherwise, instruction-tuning deteriorates the correlations over the base variants. Also, instruction-tuning fails to improve the correlations over more basic models such as GPT2-base.
@@ 5.4 Cloze-surprisal
In Table 4, we report the percentage gains in rank-biserial correlations for cloze-surprisal over direct-surprisal from same model variants across the four datasets. We find cloze-surprisal to be boosting the correlations with extra points in many cases; however, it also deteriorates the correlations heavily in other cases. For the GPT2 model family, cloze-surprisal consistently improves the corpus-based datasets’ correlations, yielding the strongest correlations in this study. Out of the 32 recorded surprisals on each dataset, cloze surprisal from GPT2-base achieves the highest correlations on VUA-ratings, with $r=.499$, $\rho=.499$, $r_{b}=.687$ and $auc=.843$. However, cloze-surprisal of GPT2 models consistently deteriorates the correlations on the synthetic datasets. Llama and Qwen model families also show positive gains from cloze surprisal in most of the cases, however, not as consistently as for GPT2 models on VUA. Most importantly, unlike the GPT2 model family, Llama and Qwen boost correlations on the synthetic datasets except for a very few cases. Cloze-surprisal achieves the overall strongest correlation on the Lai2009 dataset using the Qwen-14B model, with $r_{b}=.539$ and $auc=.779$. Also, it achieves the overall strongest correlation on the GPT-4o-metaphors dataset using the Llama-8B model, with $r_{b}=.684$ and $auc=.842$.
To investigate the effect of incorporating the whole sentence context in surprisal computation, we plot the distribution of LMs’ direct and cloze surprisals of metaphoric words in VUA in Figure 3 (Appendix A.6). We observe that the distribution of direct-surprisal is often right-skewed with many values near zero. While cloze-surprisal shifts the distribution away from zero and approaches a normal distribution. Furthermore, we plot the distribution of GPT2-base direct and cloze surprisal of metaphoric words in VUA at different sentence positions in Figure 4 (Appendix A.6). Again, we observe that cloze-surprisal shifts the distributions away from near zero values. However, against our intuition, the effect of cloze-surprisal is not more impactful in cases where the metaphor is located earlier in the sentence than later. The effect of cloze-surprisal, although positive, is not related to the position of the metaphoric word in the sentence.
@@
ppl. $r_{b}$ Nov. % $r_{b}$ Nov. %
Fiction 108 .693 2.97 .478 3.56
News 89 .653 2.80 .413 3.88
Academic 73 .588 1.85 .413 1.85
Conversation 134 .482 1.41 .800 0.62
All 96 .638 2.33 .581 2.70
@@ 6 Discussion
Across multiple novelty annotation setups—human ratings, dictionary-based binary labels, experimental contrastive items designed by experts, and LLM-generated conventional vs. novel senses—we find consistently moderate associations between LM surprisal and metaphor novelty (best: $r=.49,\rho=.50,r_{b}=.69,\mathrm{AUC}=.84$). While these results are not directly comparable to prior work due to differences in datasets and task formulations, their magnitudes are broadly in line with reported surprisal–behavior associations in acceptability judgments Tjuatja et al. (2025) and reading times Oh and Schuler (2023a). The closest point of reference in the metaphor novelty literature is Parde and Nielsen (2018), who report $r=.44$ when predicting novelty scores from a wide range of linguistic features. They additionally report moderate inter-annotator agreement, underscoring that novelty itself is challenging to measure reliably by humans. At the same time, surprisal has clear theoretical limits as a standalone predictor: interpreting a metaphor involves semantic integration and cross-domain mapping, not only predictability. We therefore expect surprisal to be most informative when combined with measures that more directly target novel interpretations and domain mappings, especially for highly creative (novel) metaphors.
@@
The negative effect of model size observed in corpus-based datasets mirrors the inverse scaling effect reported in reading time studies Oh and Schuler (2023b); Wilcox et al. (2025). Recently, Oh et al. (2024) argued that this effect is largely driven by word frequency, with larger models assigning increasingly non-human-like expectations to rare words. Since corpus-based novelty scores are often correlated with word frequency Do Dinh et al. (2018); Reimann and Scheffler (2024), frequency can be an underlying factor for this negative scaling effect in our study. By contrast, the synthetic datasets control for lexical identity and frequency: conventional and novel senses are elicited for the same set of words, and stimuli are constructed to cleanly separate the two senses. Under these controls, scaling improves alignment with novelty labels, agreeing with reports of positive scaling effects in other behavioural settings such as acceptability judgments and other reading time studies Tjuatja et al. (2025); Wilcox et al. (2023). Overall, we see our results as evidence that these diverging scaling effects are due to the nature of the dataset types. While corpus-based datasets reflect metaphor novelty mainly through lexical properties such as word frequency, synthetic datasets more directly isolate the conventional–novel distinction by controlling for lexical properties confounds.
@@ 7 Conclusion
We acknowledge a couple of limitations in this work. First, the scarcity of high-quality metaphor novelty annotations in existing literature constrains both coverage and generalizability. Second, we rely on pretrained language models whose training data and processes are not fully disclosed. Additionally, although we carefully analyse the effects of model architecture, size, and domain (genre), future work could adopt mixed-effects models to test the interaction of these variables.
@@ Acknowledgments
We acknowledge the anonymous reviewers and area chairs for their valuable comments and feedback. Furthermore, we thank Özge Alaçam, Annett Jorschick, Vicky Tzuyin Lai and Tiago Pimentel for their cooperation and responsiveness to our inquiries.
Appendix A Appendix
@@ A.1 Statistics and Examples
5. One Mr Clarke can not duck away from if he wants to avoid a second Winter of Discontent -0.094 conventional 3.165 7.454
@@ A.2 GPT-4o-metaphors Construction
A.3 Models Details
All the LMs used in this study are based on the HuggingFace HF models hub. Hence, we list here the exact HF models’ IDs that we used in our experiments. We use default HF parameters when forwarding the inputs through the model layers. For models meta-data and parameters, please refer to the models’ cards on HF using the links below:
A.4 Use of AI Assistants
AI assistants were used during manuscript preparation only for limited linguistic editing to improve clarity and style, and for writing auxiliary code (e.g., for visualisations). They were not used for scientific reasoning, evaluation decisions, or interpretation of results; all analyses and conclusions were drawn by the authors.
A.5 Scientific Artifacts
VUA datasets are publicly available datasets and intended for scientific research, and we follow this purpose in this study. Lai2009 dataset has copyright regulations; the author of this dataset approved our usage of the dataset, and she promised to share the data with anyone interested in reproducing our experiments. We double-checked that the datasets and prompts do not contain any personal data.
A.6 Distributions Visualisations
@@
Figure 3: Histograms of VUA metaphor novelty annotations, and GPT2-base direct surprisal (dark colour) of content metaphoric words (15,155) across 10 LMs. Additionally, cloze-surprisal of the same models is plotted in (dashed) light colour, and surprisal from instruct-tuned variants (whenever applicable) is plotted with borderlines.
Figure 4: Scatter plots between metaphor novelty scores of VUA-ratings on the x-axis; and direct surprisal (blue) and cloze surprisal (red) (from GPT2-base) on the y-axis at different relative positions of metaphoric words in sentences. 0-10% means metaphoric words located at the first 1/10Length1/10*Length of the sentence.